5. The Centers for Disease Control have determined that when a person is given a vaccine, the probability that the person will develop immunity to a virus is 0.8. If 7 people are given this vaccine, find the probability that
a. at least five will develop immunity



b. exactly four will develop immunity



c. all will develop immunity

Answer by edjones(8007)

Expand the binomial (a+b)^7, Let a=.8 and b=.2
.
c. .8^7=.2097
.
b. 35*(.8)^4*(.2)^3=.1147
.
a. a^7+7a^6b+21a^5b^2=.852
.
Ed